A = 1/2h(B+b)
area of a trapezoid
solve for b

Answer by sarah_adam(201)

area of a trapezoid : A = 1/2 * h *(B+b) where B and b are the lengths of the two parallel bases of a trapezoid, and h is its height
A = 1/2 * h *(B+b)
2A/h = B + b
Therefore b = 2A/h - B